OpenStudy (anonymous):

Simple Cuts is a hair salon that offers two types of haircuts: a buzz cut and a flat-top. Jerry charges $9 for a buzz cut and $13 for a flat-top. If Jerry made $600 one day and knows that he cut three times as many buzzes as flat-tops, how many flat-tops did he cut that day? A. 15 B. 21 C. 40 D. 280

OpenStudy (anonymous):

15

OpenStudy (jdoe0001):

b = buzzes cuts quantity f = flat-top cuts quantity so he charges $9 per buzz cut, so the amount he made that day for buzz cuts was 9*b = 9b so he charges $13 per flat-top cuts, so the amount he made that day for flat-top cuts was 13 * f = 13f so we can say that, the total amount, $600, is the sum of both of them 9b + 13f = 600 he made $600, but he only recalls that buzz cuts he made, were 3 times as much as the float-tops thus if he cut " f amount" of flat-tops that day, then he cut "3 times that amount", that is 3f of buzz cuts, thus b = 3f so we could say that \(\bf 9b + 13f = 600\implies 9(3f) + 13f = 600\) solve for "f" to find out how many flat-top cuts he made

OpenStudy (anonymous):

So, 13 dollars for a haircut and 9 dollars for a haircut. We don't know how many haircuts for either, so let's let the multiplier of both of those be some unknown variable, x. We do know, however, that the buzz cuts were purchased 3 times more than the flat tops, so it's 3 times the multiplier - or 3x. All set equal to the total sales of the day. Thus, 13x + 9(3x) = 600 13x + 27x = 600 40x = 600 x = 15

OpenStudy (anonymous):

Thank you guys all so much it was 15 :)